# nonblind {nonblind}

### v1

@@ Abstract
Algorithms frequently assist, rather than replace, human decision-makers. However, the design and analysis of algorithms often focus on predicting outcomes and do not explicitly model their effect on human decisions. This discrepancy between the design and role of algorithmic assistants becomes of particular concern in light of empirical evidence that suggests that algorithmic assistants again and again fail to improve human decisions. In this article, we formalize the design of recommendation algorithms that assist human decision-makers without making restrictive ex-ante assumptions about how recommendations affect decisions. We formulate an algorithmic-design problem that leverages the potential-outcomes framework from causal inference to model the effect of recommendations on a human decision-maker’s binary treatment choice. Within this model, we introduce a monotonicity assumption that leads to an intuitive classification of human responses to the algorithm. Under this monotonicity assumption, we can express the human’s response to algorithmic recommendations in terms of their compliance with the algorithm and the decision they would take if the algorithm sends no recommendation. We showcase the utility of our framework using an online experiment that simulates a hiring task. We argue that our approach explains the relative performance of different recommendation algorithms in the experiment, and can help design solutions that realize human–AI complementarity.
00footnotetext: Bryce McLaughlin (brycem@stanford.edu) and Jann Spiess (jspiess@stanford.edu), Graduate School of Business, Stanford University. We thank Sanmi Koyejo, Jason Weitze, Emma Rockall, Adrienne Propp, Carl Meyer, Omer Shiran-Cohen, and audience members at the 2023 INFORMS Annual Meeting and at Stanford for helpful comments and suggestions.
@@ 1 Introduction
Algorithms often assist, rather than replace, human decision-makers. For example, in the United States, federal pretrial service officers set bail for defendants after observing a Pretrial Risk Assessment (PTRA), rather than algorithmically implementing a bail decision in response to the PTRA. Similarly, doctors may be assisted by clinical decision support systems throughout the patient care process (Sutton et al., 2020) and managers may be assisted by pre-employment assessments throughout the hiring process (Raghavan et al., 2020). In all of these cases, the human maintains the final decision authority. However, the design and analysis of these algorithms often focus on the prediction of outcomes as if algorithmic recommendations were implemented directly, rather than how these recommendations affect human decisions.
@@
The narrow focus of existing approaches on predicting outcomes rather than optimizing for human–AI complementarity becomes of particular concern in light of a growing body of evidence that shows that human decision-makers with access to algorithmic recommendations often make worse decisions than if these recommendations had been implemented directly, and in some cases may even take worse decisions than they would have taken without access to the algorithm. In a recent review, Hemmer et al. (2021) identifies 53 articles that report (i) unassisted human performance, (ii) algorithm performance, and (iii) algorithm-assisted human performance on a decision task. Only in 16 of these does the algorithm-assisted human decision-maker perform best. In many of these studies, decision-makers exhibit noisy and biased adoption of the algorithm’s recommendations, which counteracts the gains associated with the additional information the algorithm provides. This suggests the necessity to incorporate decision-makers’ adoption decisions explicitly when designing the algorithm.
@@
This article formalizes the design of recommendation algorithms that assist humans in making binary treatment choices via an optimization problem that leverages the potential outcomes framework from causal inference. For the analysis of recommendation algorithms, we employ the notion of “potential outcomes” and make connections to the analysis of instrumental variables in causal inference with imperfect compliance. This view suggests that an optimal recommendation algorithm should take into account the final outcome of interest to which the decision is related as well as the decision-maker’s choice of when to comply with a recommendation. We then show that this framework is helpful in designing better recommendation algorithms and making sense of their empirical performance using an online experiment.
@@
In Section 2, we introduce a principal–agent model that formalizes the design of recommendation algorithms that assist a human in making binary treatment choices. As a motivating example, we consider a manager making hiring decisions on applicants who are evaluated by a pre-employment assessment. The principal in our problem is the company, who designs an algorithm that assists a hiring manager (the agent) in making the ultimate hiring decision. This model allows for complementarity in information and expertise between the principal’s algorithm and the agent, thus moving beyond models in which perfect adherence to the algorithmic recommendation is optimal.
@@
Section 3 connects our problem setup to the potential-outcomes framework from causal inference. There, we introduce a monotonicity assumption that allows us to describe the effect of recommendations in terms of compliance, similar to the analysis of instrumental variables by Imbens and Angrist (1994). The potential-outcomes framework formalizes how individual outcomes or actions change in response to treatment choices. We use this framework to capture how different recommendations given by the algorithm lead to different decisions by the human decision-maker. Using our potential outcomes, we can state the optimal recommendation algorithm as minimizing an objective that separates into two parts, (1) the loss incurred by a direct implementation of the algorithm’s recommendations and (2) the loss incurred (or avoided) due to how decision-makers deviate from the recommendations. Our monotonicity assumption then implies that recommendations can only move the decision-maker toward the recommended action.
@@
In Section 4, we showcase our framework by applying it to an online experiment that involves a hypothetical hiring task. This experiment asks subjects to make 25 hiring decisions on hypothetical applicants, and randomly varies the recommendation algorithm across subjects. This setup allows us to evaluate how the structure of recommendations impacts ultimate task performance. Specifically, the separate recommendation algorithms are based on different assumptions about compliance behaviors. In line with our framework, we find that subjects perform better when the recommendations are designed to provide complementary information, rather than optimized for hiring decisions directly. The best-performing algorithm assumes that human decision-makers comply with the algorithm only selectively in those cases where they are uncertain. The resulting decisions outperform unassisted human decisions as well as the best performance by an algorithm on its own. We thus demonstrate that our framework can help design systems that improve decisions through human–AI complementarity.
@@
Our work responds to an empirical literature that analyzes inefficiencies in the interactions between human decision-makers and prediction algorithms. This literature has identified cases where human decision-makers imperfectly use prediction algorithms in high-stakes decisions in criminal justice (Imai et al., 2021; Stevenson and Doleac, 2022; Angelova, Dobbie, and Yang, 2023), healthcare (Agarwal et al., 2023; Maron et al., 2020), and child protective services (Fogliato, De-Arteaga, and Chouldechova, 2022; Mills and Grimon, 2022), so much so that the assisted human decisions are often no better than unassisted human decisions. A more extensive review of this literature is detailed by Lai et al. (2021). Of the works in this literature, our empirical approach is most related to that of Green and Chen (2019) who posit ways human decision-makers should respond to algorithms and test these hypotheses in the lab. Relative to this literature, we focus on designing complementary recommendations, rather than focusing on describing ways in which humans misuse algorithms.
@@
We join a collection of works spanning the areas of operations, economics, and computer science that design algorithms to assist humans in taking better decisions. Bansal et al. (2019) and Donahue, Chouldechova, and Kenthapadi (2022) propose improvements to collaborative decision processes based on the private information of each agent. Balakrishnan, Ferreira, and Tong (2022) and Orfanoudaki et al. (2022) find that humans often incorrectly assume linear relationships, then develop strategies to help human–algorithm decision systems overcome this issue. Athey, Bryan, and Gans (2020) and McLaughlin and Spiess (2022) model how effort costs and reference points, respectively, affect a human’s adoption of algorithmic assistance. Baek et al. (2023) and Bastani, Bastani, and Sinchaisri (2022) apply ideas from reinforcement learning to incorporate past human actions into the recommendations of assistive algorithms. Caro and de Tejada Cuenca (2023) and Sun et al. (2022) use data on workers’ deviations from algorithm advice to increase compliance. Noti and Chen (2023) focuses on the question of when to provide assistance, and shows that strategically restricting advice can preserve complementary between machine and human performance. Our approach differs from these prior works as it provides a causal framework to capture how a human’s response varies with the underlying algorithm. Similar to the work of Ibrahim, Kim, and Tong (2021) (where human predictors assist algorithmic decision-makers), our results highlight the importance of private information in designing impactful collaboration.
@@
Our methods build upon a literature that optimizes the design of estimators that inform decision-making. We combine models of communication between heterogeneous agents (Kamenica and Gentzkow, 2011; Andrews and Shapiro, 2021; Spiess, 2022) with approaches from causal inference with imperfect compliance (Angrist, Imbens, and Rubin, 1996; Vytlacil, 2002; Mogstad and Torgovitsky, 2018) in order to design optimal recommendation algorithms in the prescriptive sense (Bertsimas and Kallus, 2020). This approach parallels work that transfers concepts from causal inference for use in decision-making (Manski, 2021, 2023), in particular the approach in Saghafian (2023) of incorporating unobserved confounding using sets of causal models and the approach in Ben-Michael et al. (2022) of learning robust decision policies from non-stochastic treatment assignments.
@@ 2 Problem Setup
We introduce a principal–agent model of joint human–algorithm decision-making{nonblind}, which builds upon McLaughlin and Spiess (2022). In this model, the principal designs an algorithm that gives recommendations to an agent. The agent is a human decision-maker who observes each recommendation and subsequently makes a binary treatment choice. As our main illustration, we apply this model to the problem a company (principal) faces when implementing an algorithmic pre-employment assessment to assist managers (agents) in the hiring process. The company wants its managers to hire good workers and not hire bad workers. The company impacts the managers’ decisions only through the recommendation algorithm they implement. We allow for the managers to have both unknown information and preferences, allowing for complex relationships between the algorithm the company implements and the decisions that managers take.
@@ 2.1 Hiring as a Binary Treatment Choice Problem
We consider the problem of taking a binary decision $D$ before a binary outcome $Y$ is realized. We capture the impact of this decision through a loss function $\ell$, which describes regret from implementing $D$ compared to the optimal decision. For each problem instance, we assume access to characteristics $X$ that have predictive power for $Y$.
As a running example throughout this paper, as the binary treatment choice we consider a hiring decision $D\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$ with the outcome $Y\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}}\}$ measuring the applicant’s job performance, which is unknown at the time of the decision. We measure the utility associated with this decision via the loss function,
@@
The type-I error cost $c_{I}>0$ captures the value of a good ( G) worker’s productivity minus their hiring cost while the type-II error cost $c_{II}>0$ captures the hiring cost minus whatever productivity a bad ( B) worker provides.
While worker performance $Y$ is ex-ante unknown, we assume that there is information $X$ collected about the worker which is available at the time $D$ is made. We do not assume any specific structure of $X$, as the data available about applicants before hiring may include their resume, written question responses, video responses to interview questions, or the actions the worker took during a game (Li et al., 2021). We do require that the information $X$ is collected, documented, and made available prior to making a hiring decision $D$ on each worker, and for simplicity assume that $X$ has finite support.
@@ 2.2 Algorithm Design as a Principal–Agent Problem
We examine the decision $D$ through the lens of a principal–agent model with asymmetric information and potentially misaligned preferences. Specifically, we assume the decision $D$ is taken by an agent who (1) observes private information $U$, (2) does not observe characteristics $X$ directly, and (3) may not share the preferences represented by the loss function $\ell$. To improve the decisions by the agent, the principal (whose preferences are captured faithfully by $\ell$) implements an algorithm $f^{\textnormal{rec}}$ that provides recommendations $R=f^{\textnormal{rec}}(X)$ to the agent. Thus, similar to the information design literature (Kamenica and Gentzkow, 2011; Bergemann and Morris, 2019), the principal has partial control over the information the agent receives before they make the final decision.
@@
Returning to our hiring example, we assume the principal is a company that implements an algorithmic recommendation system $f^{\textnormal{rec}}$ that uses the data $X$ on workers to generate hiring recommendations $R=f^{\textnormal{rec}}(X)$ which influence the hiring decisions of the agent, who is a manager at the company. The recommendations reduce the unstructured applicant data $X$ to a simpler form that is still informative about the worker’s ability $Y$. For simplicity, we limit the algorithm to either recommend a decision to the manager (not hire, $R=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$, or hire, $R=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}$) or not make a recommendation at all ($R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$). In practice, recommendations may present more granular predictive information, such as a score that represents a scaled probability a worker is good ($Y=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}}$) and allows the managers to understand the algorithm’s relative ranking of different applicant workers. It may also include an explanation of the score, highlight pieces of $X$ the algorithm deems particularly relevant, or include certainty information about the confidence of its prediction (Raghavan et al., 2020).
The manager responds to the recommendation by making a hiring decision $D$ which may depend on the recommendation $R$ along with their private information $U$ and unknown private preferences. The manager may receive this private information $U$ as a result of interviews with the applicant and may have preferences not captured by $\ell$ which drive their hiring decisions. We do not, for now, assume any specific structure of the managers’ preferences or private information. Furthermore, similar to how we assume the company does not observe $U$, we assume the manager does not observe $X$. However, we allow arbitrary correlation between $X$ and $U$, effectively capturing the idea that $X$ and $U$ contain shared information. This flexibility allows us to capture the complex decisions managers exhibit empirically. Note that $D$ generally depends on the recommendation rule $f^{\textnormal{rec}}$ beyond the actual recommendation $R$ that is given in a specific case. This is because interpreting the recommendation $R$ requires an understanding of which cases lead to each recommendation.
@@
Throughout, we include the option that the algorithm does not make a recommendation, $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$, forcing the manager to make an active choice. We believe that this design feature has formal advantages for modeling recommendations and practical relevance for achieving human–algorithm complementarity. First, allowing for instances without recommendations enables us to include decisions taken without any recommendation algorithm (i.e., unassisted decisions) as a special case in our framework. Second, {nonblind} building upon McLaughlin and Spiess (2022), forcing active choices by avoiding recommendations in specific cases may reduce distortions in cases where human decision-makers know better. These beliefs are further supported by the work of Noti and Chen (2023), which finds that partially withholding algorithmic assistance (i) increases humans’ reliance on advice when given and (ii) can be leveraged strategically to improve humans’ prediction accuracy.

### v2

@@ Abstract
Algorithms frequently assist, rather than replace, human decision-makers. However, the design and analysis of algorithms often focus on predicting outcomes and do not explicitly model their effect on human decisions. This discrepancy between the design and role of algorithmic assistants becomes particularly concerning in light of empirical evidence that suggests that algorithmic assistants again and again fail to improve human decisions. In this article, we formalize the design of recommendation algorithms that assist human decision-makers without making restrictive ex-ante assumptions about how recommendations affect decisions. We formulate an algorithmic-design problem that leverages the potential-outcomes framework from causal inference to model the effect of recommendations on a human decision-maker’s binary treatment choice. Within this model, we introduce a monotonicity assumption that leads to an intuitive classification of human responses to the algorithm. Under this assumption, we can express the human’s response to algorithmic recommendations in terms of their compliance with the algorithm and the active decision they would take if the algorithm sends no recommendation. We showcase the utility of our framework using an online experiment that simulates a hiring task. We argue that our approach can make sense of the relative performance of different recommendation algorithms in the experiment and can help design solutions that realize human–AI complementarity. Finally, we leverage our approach to derive minimax optimal recommendation algorithms that can be implemented with machine learning using limited training data.
00footnotetext: Bryce McLaughlin (bryce.h.mclaughlin@gmail.com), Wharton Healthcare Analytics Lab, University of Pennsylvania and Jann Spiess (jspiess@stanford.edu), Graduate School of Business, Stanford University. We thank Maya Balakrishnan, Sanmi Koyejo, Jason Weitze, Emma Rockall, Adrienne Propp, Carl Meyer, Omer Shiran-Cohen, Haosen Ge, and audience members at the 2023 & 2024 INFORMS Annual Meetings, 2024 MSOM Conference, Stanford, and Wharton for helpful comments and suggestions.
@@ 1 Introduction
Algorithms often assist, rather than replace, human decision-makers. For example, in the United States, federal pretrial service officers set bail for defendants after observing a pretrial risk assessment (Cadigan and Lowenkamp, 2011), rather than letting the algorithm take a bail decision automatically. Similarly, doctors may be assisted by clinical decision support systems throughout the patient care process (Sutton et al., 2020) and managers may be assisted by pre-employment assessments throughout the hiring process (Raghavan et al., 2020). In all of these cases, the human maintains the final decision authority. However, the design and analysis of these algorithms often focus on the prediction of outcomes as if algorithmic recommendations were implemented directly, rather than on how these recommendations affect human decisions.
@@
The narrow focus of existing approaches on predicting outcomes rather than optimizing for human–AI complementarity becomes of particular concern in light of a growing body of evidence that shows that human decision-makers with access to algorithmic recommendations often make worse decisions than if these recommendations had been implemented directly and, in some cases, may even make worse decisions than they would have taken without access to the algorithm. Hemmer et al. (2021) identifies 53 articles that report (i) unassisted human performance, (ii) algorithm performance, and (iii) algorithm-assisted human performance on a decision task. Only in 16 of these does the algorithm-assisted human decision-maker perform best. In many of these studies, decision-makers exhibit noisy and biased adoption of the algorithm’s recommendations, which counteracts the gains associated with the additional information the algorithm provides. This suggests the necessity to incorporate decision-makers’ adoption decisions explicitly when designing the algorithm.
@@
This article formalizes the design of recommendation algorithms that assist humans in making binary treatment choices via an optimization problem that leverages the potential-outcomes framework from causal inference. To analyze recommendation algorithms, we employ the notion of “potential decisions” and make connections to the analysis of instrumental variables in causal inference with imperfect compliance. This view suggests that an optimal recommendation algorithm should take into account the final outcome of interest to which the decision is related as well as the decision-maker’s choice of when to comply with a recommendation. We then show that this framework helps design better recommendation algorithms and make sense of their empirical performance through an online experiment. Finally, we derive robust recommendation algorithms that achieve minimax performance from limited training data. While our theorems and propositions explicitly hold for this stylized setup, the framework we develop is broadly applicable for designing algorithms that assist human decision-makers. Specifically, we connect algorithm design choices to the assumptions the designer is willing to make about the behavior of the human decision-maker, exposing the central role of human decision data to the design of these tools.
@@
In Section 2, we introduce a principal–agent model that formalizes the design of AI recommendation systems that assist a human in making binary treatment choices. As a motivating example, we consider a manager making hiring decisions on applicants evaluated by a pre-employment assessment. The principal in our problem is the company, which designs an algorithm that assists a hiring manager (the agent) in making the ultimate hiring decision. This model allows for complementarity in information and expertise between the principal’s algorithm and the agent, thus moving beyond models in which perfect adherence to the algorithmic recommendation is optimal.
@@
Section 3 connects our problem setup to the potential-outcomes framework from causal inference. There, we introduce a monotonicity assumption that allows us to describe the effect of recommendations in terms of compliance, similar to the analysis of instrumental variables by Imbens and Angrist (1994). The potential-outcomes framework formalizes how individual outcomes or actions change in response to treatment choices. We use this framework to capture how different recommendations given by the algorithm lead to different decisions by the human decision-maker. Using our “potential decisions”, we can state the optimal recommendation algorithm as minimizing an objective that separates into two parts, (1) the loss incurred by a direct implementation of the algorithm’s recommendations and (2) the loss incurred (or avoided) due to how decision-makers deviate from the recommendations. Our monotonicity assumption then simplifies the second part since it implies that recommendations can only move the decision-maker toward the recommended action.
@@
In Section 4, we showcase our framework by applying it to an online experiment that involves a hypothetical hiring task. This experiment asks subjects to make 25 hiring decisions and randomly varies the recommendation algorithm across subjects. This setup allows us to evaluate how the structure of recommendation algorithms impacts ultimate task performance. Specifically, the separate recommendation algorithms are based on different assumptions about compliance behaviors. In line with our framework, we find that subjects perform better when the recommendations are designed to provide complementary information, rather than optimized for hiring decisions directly. The best-performing algorithm assumes that human decision-makers comply with the algorithm selectively, only impacting the cases where the human is uncertain. The decisions assisted by this algorithm outperform unassisted human decisions as well as the best performance by an algorithm on its own. We thus demonstrate that our framework can help design systems that improve decisions through human–AI complementarity.
Finally, in Section 5, we leverage our approach to find robust recommendation algorithms from limited data. We focus on the realistic case where the principal may only have data about unassisted human decisions available and thus has to rely on assumptions about compliance and active decisions. Assuming that the decision-maker is more likely to comply in cases where they do not make good decisions themselves, we derive minimax optimal recommendation rules. Our result shows that learning successful recommendations is about learning from mistakes rather than from the relationship of outcomes to characteristics alone. We end by discussing the resulting robust optimization problem that can be implemented using machine learning.
@@
Our work responds to an empirical literature that analyzes inefficiencies in the interactions between human decision-makers and prediction algorithms. This literature has identified cases where human decision-makers imperfectly use prediction algorithms in high-stakes decisions in criminal justice (Imai et al., 2021; Stevenson and Doleac, 2022; Angelova, Dobbie, and Yang, 2023), healthcare (Agarwal et al., 2023; Maron et al., 2020), and child protective services (Fogliato, De-Arteaga, and Chouldechova, 2022; Mills and Grimon, 2022), so much so that the assisted human decisions are often no better than unassisted human decisions. A more extensive review of this literature is detailed by Lai et al. (2021). Of the works in this literature, our empirical approach is most related to that Green and Chen (2019), which posits ways human decision-makers should respond to algorithms and tests these hypotheses in the lab. Relative to this literature, we focus on designing complementary recommendations, rather than focusing on describing ways in which humans misuse algorithms.
@@
We join a collection of works spanning the areas of operations, economics, and computer science that design algorithms to assist humans in making better decisions. Bansal et al. (2019) and Donahue, Chouldechova, and Kenthapadi (2022) propose improvements to collaborative decision processes based on the private information of each agent. Balakrishnan, Ferreira, and Tong (2022) and Orfanoudaki et al. (2022) find that humans often incorrectly assume linear relationships and then develop strategies to help human–algorithm decision systems overcome this issue. Athey, Bryan, and Gans (2020) and McLaughlin and Spiess (2022) model how effort costs and reference points, respectively, affect a human’s adoption of algorithmic assistance. Snyder, Keppler, and Leider (2023) shows how workers’ compliance with an algorithm’s recommendation in a service system varies with the system’s utilization. Baek et al. (2023) and Bastani, Bastani, and Sinchaisri (2022) apply ideas from reinforcement learning to incorporate past human actions into the recommendations of assistive algorithms. Caro and de Tejada Cuenca (2023) and Sun et al. (2022) use data on workers’ deviations from algorithm advice to increase compliance. Noti and Chen (2023) focuses on when to provide assistance, and shows that strategically restricting advice can preserve the complementary strengths of the algorithm and human. Our approach differs from these prior works as it provides a causal framework to capture how a human’s response varies with the underlying algorithm. Similar to the work of Ibrahim, Kim, and Tong (2021) (where human predictors assist algorithmic decision-makers), our results highlight the importance of private information in designing impactful collaboration.
@@
Our methods build upon a literature that optimizes the design of estimators that inform decision-making. We combine models of communication between heterogeneous agents (Kamenica and Gentzkow, 2011; Andrews and Shapiro, 2021; Spiess, 2022) with approaches from causal inference with imperfect compliance (Angrist, Imbens, and Rubin, 1996; Vytlacil, 2002; Mogstad and Torgovitsky, 2018) to design optimal recommendation algorithms in the prescriptive sense (Bertsimas and Kallus, 2020). This approach parallels work that transfers concepts from causal inference for use in decision-making (Manski, 2021, 2023), in particular the approach in Saghafian (2023) of incorporating unobserved confounding using sets of causal models and the approach in Ben-Michael et al. (2022) of learning robust decision policies from non-stochastic treatment assignments.
@@ 2 Problem Setup
We introduce a principal–agent model of joint human–algorithm decision-making{nonblind}, which builds upon McLaughlin and Spiess (2022). In this model, the principal designs an AI system that can give recommendations to an agent. The agent is a human decision-maker who observes each recommendation and subsequently makes a binary treatment choice. As our main illustration, we apply this model to the problem a company (principal) faces when implementing an algorithmic pre-employment assessment to assist a manager (agent) in the hiring process. The company wants its managers to hire good workers and not hire bad workers. The company impacts the managers’ decisions only through the recommendation algorithm they implement. We allow for the managers to have both unknown information and preferences, allowing for complex relationships between the algorithm the company implements and the decisions that managers take.
@@ 2.1 Hiring as a Binary Treatment Choice Problem
We consider the problem of taking a binary decision $D$, after which a binary outcome $Y$ is revealed. We capture the impact of this decision through a loss function $\ell$, which describes the loss from implementing $D$ compared to the optimal decision. For each problem instance, we assume access to characteristics $X$ that have predictive power for $Y$.
As a running example throughout this paper, we consider a hiring decision $D\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}\}$ as the binary treatment choice with the ability $Y\in\{\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}B}},\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}}\}$ capturing the quality of work produced by the applicant, which is unknown at the time of the decision. We measure the (dis-)utility associated with this decision via the loss function,
@@
The type-I error cost $c_{I}>0$ captures the cost to the company of not hiring a good ( G) worker (relative to hiring them). The type-II error cost $c_{II}>0$ captures the cost of hiring a bad ( B) worker (relative to not hiring them).
While worker ability $Y$ is ex-ante unknown, we assume that there is data $X$ collected about the worker that can be accessed by an algorithm ahead of the decision $D$ being made. We do not assume any specific structure of $X$, as the data available about applicants before hiring may include their resume, written question responses, video responses to interview questions, or the actions the worker took during a game (Li et al., 2021). We do require that the information $X$ is collected, documented, and made available prior to making a hiring decision $D$ on each worker, and for simplicity assume throughout that $X$ has finite and full support.
@@ 2.2 Algorithm Design as a Principal–Agent Problem
We examine the decision $D$ through the lens of a principal–agent model with asymmetric information and potentially misaligned preferences. Specifically, we assume the decision $D$ is taken by an agent who (1) observes private information $U$, (2) does not observe characteristics $X$ directly, and (3) may not share the preferences represented by the loss function $\ell$. We allow for arbitrary overlap between the information in $X$ and $U$, so the model allows for some or even all of the characteristics $X$ to be available to the agent.
To improve the decisions by the agent, the principal (whose preferences are captured faithfully by $\ell$) implements an algorithm $f^{\textnormal{rec}}$ that provides recommendations $R=f^{\textnormal{rec}}(X)$ to the agent. Thus, similar to the information design literature (Kamenica and Gentzkow, 2011; Bergemann and Morris, 2019), the principal has partial control over the information presented to the agent before the agent makes the final decision. We think of the algorithm as provided by an AI system that may learn from training data, but we abstract from the training process and focus on the resulting recommendation rule $f^{\textnormal{rec}}$ itself.
@@
Returning to our hiring example, we assume the principal is a company that implements an algorithmic recommendation system $f^{\textnormal{rec}}$ that uses the characteristics $X$ on workers to generate hiring recommendations $R=f^{\textnormal{rec}}(X)$, which influence the hiring decisions of the agent, who is a manager at the company. The recommendations reduce the unstructured applicant data $X$ to a simpler form that may still be informative about the worker’s ability $Y$. For simplicity, we limit the algorithm to either recommend a decision to the manager (not hire, $R=\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}N}}$, or hire, $R=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} H}}$) or not make a recommendation at all ($R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$). In practice, recommendations may present more granular predictive information, such as a score that represents a scaled probability that a worker is good ($Y=\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} G}}$) and allows the managers to understand the algorithm’s relative ranking of different workers. It may also include an explanation of the score, highlight pieces of $X$ the algorithm deems particularly relevant, or include certainty information about the confidence of its prediction (Raghavan et al., 2020).
The manager responds to the recommendation by making a hiring decision $D$ which may depend on the recommendation $R$ along with their private information $U$ and unknown private preferences. This information $U$ consists of all knowledge the manager has of the applicant, including potentially some of the information contained in $X$ (such as the candidate’s resume) as well as new information the manager learns through their own conversations with the applicant (such as prior relationships with employees at the company or interest in areas the company is looking to expand). Thus, we allow arbitrary correlation between $X$ and $U$, but do not allow the manager to observe $X$ directly. This effectively captures the idea that $X$ and $U$ contain shared information, but that neither necessarily dominates the other. Additionally, the manager may exhibit preferences that do not directly align with $\ell$, preferring to hire individuals with whom they get along well (or have prior relationships) even when those applicants are less likely to be good than others. This flexibility allows us to capture the complex decisions managers exhibit empirically. We summarize the relationship between the elements of our model in Figure 1.
$X$$U$$D$$R$$Y$$\ell(Y,D)$$f^{\textnormal{rec}}$
Figure 1: This diagram illustrates how the elements of our model are related to each other. The algorithm’s information $X$ determines the recommendation $R$ through the algorithm $f^{\textnormal{rec}}$. This recommendation in turn impacts the agent’s (manager’s) decisions $D$ in a way that depends on $f^{\textnormal{rec}}$. The agent’s decision is also impacted by their information $U$. Both the algorithm’s information $X$ and the agent’s information $U$ can be predictive of the outcome (ability) $Y$. This outcome, combined with the agent’s decision, determines the loss $\ell(Y,D)$. The dashed line between $X$ and $U$ indicates that they may contain common information. The shading of $U$ illustrates the challenge that, from the perspective of the principal (firm), the private information of the agent (manager) may never be directly observed.
@@
Throughout, we include the option that the algorithm does not make a recommendation, $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$, forcing the manager to make an active decision. We believe that this design feature has formal advantages for modeling recommendations and practical relevance for achieving human–algorithm complementarity. First, we can consider unassisted decisions, those taken without any recommendation algorithm, as equivalent to those taken under a recommendation algorithm which always sends $R={{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\varnothing}}$, making unassisted decisions a special case of our framework. Second, {nonblind} building upon McLaughlin and Spiess (2022), forcing active decisions by avoiding recommendations in specific cases may reduce distortions in cases where human decision-makers know better. These beliefs are further supported by the work of Noti and Chen (2023), which finds that partially withholding algorithmic assistance (i) increases humans’ reliance on advice when given and (ii) can be leveraged strategically to improve humans’ prediction accuracy.